# General targeted machine learning for modern causal mediation analysis

### v1

@@ 4 Estimators and their sampling distribution
Having derived the corresponding von Mises expansions, in this section we focus on the development of estimation algorithms. These estimators have multiple ingredients. First, it is necessary to specify an algorithm for the construction of a plug-in estimator. This estimator will be constructed by a sequential regression procedure using the definitions of $\theta_{j}$ and $b_{j}$ given in expressions (4) and (6). For instance, to construct a plug-in estimator for $\psi^{N}$, we first regress the outcome $Y_{i}$ on $(A_{i},Z_{i},M_{i},W_{i})$, using flexible data-adaptive regression procedures to obtain an estimate $\hat{\theta}_{3}^{N}$. We then use this estimate to compute the pseudo-outcome $b_{3}^{N}(a_{1},Z_{i},M_{i},W_{i};\hat{\theta}_{3}^{N})$ by computing the predictions of $\hat{\theta}_{3}^{N}$ in a new dataset where $A_{i}$ is fixed to $a_{1}$. This pseudo-outcome is then regressed on $(A_{i},Z_{i},W_{i})$ to obtain an estimate $\hat{\theta}_{2}^{N}$, which is then used to construct a new pseudo-outcome $b_{2}^{N}(a_{2},Z_{i},W_{i};\hat{\theta}_{2}^{N})$. This pseudo-outcome is regressed on $(A_{i},W_{i})$, leading to an estimate $\hat{\theta}_{1}^{N}$, which is used to compute $b_{1}^{N}(W_{i},\hat{\theta}_{1}^{N})$. The plug-in estimator for $\psi^{N}$ is an average across $i$ of $b_{1}^{N}(W_{i},\hat{\theta}_{1}^{N})$. By the analyses of the previous section, these plug-in estimators will have a first order bias of magnitude $-\mathsf{E}[\varphi(X,\hat{\eta}^{N})]$. As described in the previous section, an estimate of this bias can be added back to the plug-in estimator to obtain a “de-biased” one-step estimator. In general, while the expectation in $\mathsf{E}[\varphi(X,\hat{\eta}^{E})]$ can be estimated with an empirical average, and the parameters $\theta^{E}$ in $\eta^{E}$ can be estimated with sequential regression, it remains to develop estimators for the parameters $\alpha^{E}$ in expressions (5) and (7). If $M$ and $Z$ are categorical, these parameters may be estimated by plugging in probability estimates obtained using off-the-shelf machine learning classification methods. However, if either of these variables is continuous or high-dimensional, plug-in estimation becomes computationally intractable, as it would require estimating conditional densities and/or integrals with respect to those densities. To avoid this problem, we instead adopt ideas from a recent strand of articles on Riesz learning (Chernozhukov et al., 2022a, b).
@@
The foundational idea behind Riesz learning is that for an arbitrary $b$, if $\mathsf{E}[b(X,\theta)]$ is a continuous linear functional of $\theta$, then by the Riesz representation theorem there exists a unique function $\alpha$ such that $\mathsf{E}[b(X;\theta)]=\mathsf{E}[\alpha(X)\theta(X)]$ for all $\theta$ with $\mathsf{E}[\theta(X)^{2}]<\infty$. In the proof of the proposition below we show that the parameters in (5) and (7) correspond to Riesz representers of functionals that involve the $b$ functions in (4) and (6). The mean squared error (MSE) loss function then provides an approach to estimate the $\alpha$ parameter directly, avoiding estimation of densities and integrals required for the plug in estimator. Specifically, we have
@@ 5 Extension to non-binary treatments using modified treatment policies
The developments in the previous sections assumed that the exposure variable $A$ is binary, and that interest lies in decomposing the average treatment effect, defined as a contrast between outcome expectations in a hypothetical world where $\mathsf{P}(A=1)=1$ vs. a hypothetical world where $\mathsf{P}(A=0)=1$. In this section we focus on an extension of the methods to problems where the exposure is non-binary, and where the focus is on decomposing effects other than the average treatment effect. In particular, we focus on the methodology of modified treatment policies (MTP, Díaz and van der Laan, 2012; Haneuse and Rotnitzky, 2013; Díaz et al., 2023). Let $d_{1}(a,w)$ and $d_{0}(a,w)$ denote two functions that map a treatment value $a$ amd a covariate value $w$ into potential post-intervention treatment values. Causal effects of modified treatment policies are defined as contrasts $\mathsf{E}[Y(d_{1})-Y(d_{0})]$, where $Y(d)$ denotes the counterfactual outcome that would have been observed in a hypothetical world where treatment had been assigned according to $d(A,W)$, where we note that the MTP definition recovers the average treatment effect of a binary exposure by setting $d_{1}=1$ and $d_{0}=0$. Modified treatment policies allow the user to define interesting causal contrasts for continuous or multivariate exposures by careful consideration of the functions $d$ being contrasted. For instance, a researcher interested in the effect of an increase of 10% in a continuous exposure might consider setting
@@
and $d_{0}(A,W)=A$, where $u(w)$ is largest feasible value of $A$ within strata $W=w$. For a complete discussion on the flexibility of the MTP framework and other interesting policy definitions we refer the reader to the original articles as well as the reviews by Hoffman et al. (2023) and Williams and Díaz (2023).
@@ 6 Numerical studies
Here, $(\epsilon,\lambda_{1},\lambda_{2},\gamma_{1},\gamma_{2})^{\top}=(0.5,0.4,0.6,0.6,0.4)^{\top}$ are pre-specified parameters; $Be(\alpha,\beta)$ denotes a Beta distribution with shape parameters $\alpha$ and $\beta$; $Bern(\pi)$ denotes a Bernoulli distribution with probability $\pi$; $N(\mu,\sigma^{2})$ denotes a normal distribution with mean $\mu$ and variance $\sigma^{2}$; $TN(a,b,\mu,\sigma^{2})$ denotes a truncated normal distribution derived from $N(\mu,\sigma^{2})$, but with values in $[a,b]$.
@@
For estimating the Riesz representers $\alpha$ we use deep learning (LeCun et al., 2015) using the loss functions in (8) and (9). For the sequential regression function $\theta$, we use an ensemble of LightGBM (Ke et al., 2017), mean function, Multivariate Adaptive Regression Splines (Friedman, 1991) and neural networks, based on the Super Learner algorithm (van der Laan et al., 2007). The simulation study is conducted by drawing 500 datasets of sizes $n\in\{500,1000,2000\}$. The R code for implementation is publicly available on Github. We note that this simulation setup is representative of typical practice in statistics, and that we have no reason to expect that both the Riesz representers and the sequential regressions estimators to not be consistent. Therefore, the properties of Theorem 2 are not expected to hold exactly.
@@ Appendix B Auxiliary results
Appendix C Proofs of results in the main manuscript
C.1 Theorem 1

### v2

@@ 4 Estimators and their sampling distribution
Having derived the corresponding von Mises expansions, in this section we focus on the development of estimation algorithms. These estimators have multiple ingredients. First, it is necessary to specify an algorithm for the construction of a plug-in estimator. This estimator will be constructed by a sequential regression procedure using the definitions of $\theta_{j}$ and $b_{j}$ given in expressions (4) and (6). For instance, to construct a plug-in estimator for $\psi^{N}$, we first regress the outcome $Y_{i}$ on $(A_{i},Z_{i},M_{i},W_{i})$, using flexible data-adaptive regression procedures to obtain an estimate $\hat{\theta}_{3}^{N}$. We then use this estimate to compute the pseudo-outcome $b_{3}^{N}(a_{1},Z_{i},M_{i},W_{i};\hat{\theta}_{3}^{N})$ by computing the predictions of $\hat{\theta}_{3}^{N}$ in a new dataset where $A_{i}$ is fixed to $a_{1}$. This pseudo-outcome is then regressed on $(A_{i},Z_{i},W_{i})$ to obtain an estimate $\hat{\theta}_{2}^{N}$, which is then used to construct a new pseudo-outcome $b_{2}^{N}(a_{2},Z_{i},W_{i};\hat{\theta}_{2}^{N})$. This pseudo-outcome is regressed on $(A_{i},W_{i})$, leading to an estimate $\hat{\theta}_{1}^{N}$, which is used to compute $b_{1}^{N}(W_{i},\hat{\theta}_{1}^{N})$. The plug-in estimator for $\psi^{N}$ is an average across $i$ of $b_{1}^{N}(W_{i},\hat{\theta}_{1}^{N})$. By the analyses of the previous section, these plug-in estimators will have a first order bias of magnitude $-\mathsf{E}[\varphi(X,\hat{\eta}^{N})]$. As described in the previous section, an estimate of this bias can be added back to the plug-in estimator to obtain a “de-biased” one-step estimator. In general, while the expectation in $\mathsf{E}[\varphi(X,\hat{\eta}^{E})]$ can be estimated with an empirical average, and the parameters $\theta^{E}$ in $\eta^{E}$ can be estimated with sequential regression, it remains to develop estimators for the parameters $\alpha^{E}$ in expressions (5) and (7). If $M$ and $Z$ are categorical, these parameters may be estimated by plugging in probability estimates obtained using off-the-shelf machine learning classification methods. However, if either of these variables is continuous or high-dimensional, plug-in estimation becomes computationally intractable, as it would require estimating conditional densities and/or integrals with respect to those densities.
@@
To avoid this problem, we instead adopt ideas from a recent strand of articles on Riesz learning (chernuzhukov2022riesznet; Chernozhukov et al., 2022). The foundational idea behind Riesz learning is that for an arbitrary $b$, if $\mathsf{E}[b(X,\theta)]$ is a continuous linear functional of $\theta$, then by the Riesz representation theorem there exists a unique function $\alpha$ such that $\mathsf{E}[b(X;\theta)]=\mathsf{E}[\alpha(X)\theta(X)]$ for all $\theta$ with $\mathsf{E}[\theta(X)^{2}]<\infty$. In the proof of the proposition below we show that the parameters in (5) and (7) correspond to Riesz representers of functionals that involve the $b$ functions in (4) and (6). The mean squared error (MSE) loss function then provides an approach to estimate the $\alpha$ parameter directly, avoiding estimation of densities and integrals required for the plug in estimator. Specifically, we have
@@
Remark 2.
Prior literature on the development of doubly-robust estimators for mediational effects used a reparameterization of the Riesz representers for estimation (e.g, Tchetgen Tchetgen, 2013; Zheng and van der Laan, 2012; Díaz et al., 2021a). For example, in $\alpha_{2}^{N}$, by noting that $\frac{\mathsf{P}(z\mid a_{3},w)}{\mathsf{P}(z\mid a,w)}=\frac{\mathsf{P}(a_{3}\mid z,w)\mathsf{P}(a\mid w)}{\mathsf{P}(a\mid z,w)\mathsf{P}(a_{3}\mid w)}$, one could transform estimation of the density ratio with respect to $Z$ to estimation of density ratios with respect to the exposure $A$. However, in general, the assumptions required to analyze the sampling distribution of such an estimator become more difficult to satisfy. For further details, we refer the reader to Assumption 1 in §4.2 of Rudolph et al. (2024a).
@@ 5 Extension to non-binary treatments using modified treatment policies
The developments in the previous sections assumed that the exposure variable $A$ is binary, and that interest lies in decomposing effects defined as a contrast between outcome expectations in a hypothetical world where $\mathsf{P}(A=1)=1$ vs. a hypothetical world where $\mathsf{P}(A=0)=1$. In this section we focus on an extension of the methods to problems where the exposure is non-binary, and where the focus is on decomposing effects other than the average treatment effect. In particular, we focus on the methodology of modified treatment policies (MTP, Díaz and van der Laan, 2012; Haneuse and Rotnitzky, 2013; Díaz et al., 2023). MTPs have also been called dynamic interventions the depend on the natural value of treatment (Robins et al., 2004; Young et al., 2014; Richardson and Robins, 2013). Let $d_{1}(a,w)$ and $d_{0}(a,w)$ denote two functions that map a treatment value $a$ amd a covariate value $w$ into potential post-intervention treatment values. Causal effects of modified treatment policies are defined as contrasts $\mathsf{E}[Y(d_{1})-Y(d_{0})]$, where $Y(d)$ denotes the counterfactual outcome that would have been observed in a hypothetical world where treatment had been assigned according to $d(A,W)$, where we note that the MTP definition recovers the average treatment effect of a binary exposure by setting $d_{1}=1$ and $d_{0}=0$. Modified treatment policies allow the user to define interesting causal contrasts for continuous or multivariate exposures by careful consideration of the functions $d$ being contrasted. For instance, a researcher interested in the effect of an increase of 10% in a continuous exposure might consider setting
@@
and $d_{0}(A,W)=A$, where $u(w)$ is largest feasible value of $A$ within strata $W=w$. For a discussion on the flexibility of the MTP framework and other interesting policy definitions we refer the reader to the original articles as well as the review by Hoffman et al. (2023).
@@ 6 Numerical studies
Here, $W=(W_{1},W_{2},W_{3})^{\top}$ is a three-dimensional random vector.
$(\epsilon,\lambda_{1},\lambda_{2},\gamma_{1},\gamma_{2})^{\top}=(0.5,0.4,0.6,0.6,0.4)^{\top}$ are pre-specified parameters; $Be(\alpha,\beta)$ denotes a Beta distribution with shape parameters $\alpha$ and $\beta$; $Bern(\pi)$ denotes a Bernoulli distribution with probability $\pi$; $N(\mu,\sigma^{2})$ denotes a normal distribution with mean $\mu$ and variance $\sigma^{2}$; $TN(a,b,\mu,\sigma^{2})$ denotes a truncated normal distribution derived from $N(\mu,\sigma^{2})$, but with values in $[a,b]$.
@@
For estimating the Riesz representers $\alpha$ we use deep learning (LeCun et al., 2015) using the loss functions in (8) and (9). For the sequential regression function $\theta$, we use an ensemble of LightGBM (Ke et al., 2017), mean function, Multivariate Adaptive Regression Splines (Friedman, 1991) and neural networks, based on the Super Learner algorithm (van der Laan et al., 2007). The simulation study is conducted by drawing 500 datasets of sizes $n\in\{500,1000,2000\}$. The R code for implementation is publicly available on Github. We note that this simulation setup is representative of typical practice in statistics, and that we have no reason to expect that both the Riesz representers and the sequential regressions estimators are consistent. Therefore, the properties of Theorem 2 are not expected to hold exactly.
@@ 7 Illustrative Examples
8 Discussion
There has been important progress in the non-parametric definition and identification of causal parameters in mediation analysis. These parameters remain underutilized in practice because estimators for many approaches remain underdeveloped. For instance, few or no methods can handle mediators or intermediate confounders that are continuous or multivariate, a situation that is pervasive in contemporary research.
In this manuscript, we made several contributions that will increase the utilization of these parameters in practice. First, we developed non-parametric double machine learning estimators for the six mediation parameters listed in Table 1, and proved that the estimators enjoy desirable statistical properties including asymptotic linearity, weak convergence and double robustness under assumptions. To do so, we unified the identification formulas of these parameters as two statistical estimands, and then developed non-parametric estimators of these estimands with the aid of the semi-parametric efficiency theory. Our estimators are based on the EIF, which characterizes the first-order bias of plug-in estimators and non-parametric efficiency bound. The key challenge in mediation analysis with continuous or high-dimensional mediators is that estimation of the EIF requires estimation of densities of the mediator as well as integrals with respect to such densities. We address challenges in the estimation of multivariate or high-dimensional densities on the mediator by estimating $\alpha_{j}^{E},j\in\{1,2,3,4\},E\in\{N,R\}$ directly through the use of so-called Riesz learning, without the necessity of knowing their forms. We address integration with respect to the densities of the mediator by the unification, parameterizing the integrals as conditional expectations using stratified random permutations ($Z^{\pi}$) of the observed variable $Z$, which we discussed in detail in §4.3.
Second, we proposed estimation methods that can handle continuous or multivariate treatments, another pervasive situation for which methods remained underdeveloped. Third, we developed a publicly available R package crumble. The success in developing a general estimation framework for all six mediational parameters allows crumble to be general, where researchers can choose their parameters of interest based on their study questions without worrying about data complexity.
There are at least two possible future directions for this work. First, for practical reasons, we used deep learning to solve the sequential optimization problems in Proposition 2, since it is one of the few machine learning regression procedures with off-the-shelf software that allows specification of custom loss functions. The development of other optimization methods for Proposition 2, such as ensembles of regression trees or gradient boosting (Lee and Schuler, 2025), is subject to future work. Second, we note that two of the six mediational parameters quantified the path-specific effects. However, in the case of multiple mediators (e.g. two mediators $M_{1},M_{2}$), the path-specific effects can only be quantified by treating the mediators as a bundle. The definition, identification and non-parametric estimation of path-specific effects in the case of multiple mediators is another interesting future direction.
@@ Appendix B Auxiliary results
Appendix C Proofs of Proposition 1
For natural effects (Pearl, 2001), we define
$\psi^{natural}(a_{1},a_{2})=\int_{w}\int_{m}\mathsf{E}(Y\mid A=a_{1},m,w)\,\mathrm{d}\mathsf{P}(m\mid A=a_{2},w)\,\mathrm{d}\mathsf{P}(w),\$
then $\operatorname{NIE}=\psi^{natural}(1,1)-\psi^{natural}(1,0)$, and $\operatorname{NDE}=\psi^{natural}(1,0)-\psi^{natural}(0,0)$. Note that
$\psi^{natural}(a_{1},a_{2})$ $=\int_{w}\int_{m}\mathsf{E}(Y\mid A=a_{1},m,w)\,\mathrm{d}\mathsf{P}(m\mid A=a_{2},w)\,\mathrm{d}\mathsf{P}(w)$
$=\int_{w}\int_{m}\mathsf{E}(Y\mid A=a_{1},z,m,w)\,\mathrm{d}\mathsf{P}(m\mid A=a_{2},w)\,\mathrm{d}\mathsf{P}(w)\quad(\text{$Z$ is null})$
$=\int_{w}\int_{z}\int_{m}\mathsf{E}(Y\mid A=a_{1},z,m,w)\,\mathrm{d}\mathsf{P}(m\mid A=a_{2},z,w)\,\mathrm{d}\mathsf{P}(z\mid A=a_{2},w)\,\mathrm{d}\mathsf{P}(w)$
$=\mathsf{E}\bigg{[}\mathsf{E}\Big{[}\mathsf{E}\big{[}\mathsf{E}(Y\mid A=a_{1},Z,M,W)\,\big{|}\,A=a_{2},Z,W\big{]}\,\Big{|}\,A=a_{2},W\Big{]}\bigg{]},$
$=\psi^{N}(a_{1},a_{2},a_{2}).$
This shows $\operatorname{NDE}=\psi^{N}(1,0,0)-\psi^{N}(0,0,0)$ and $\operatorname{NIE}=\psi^{N}(1,1,1)-\psi^{N}(1,0,0)$.
We skip the decision theoretic effects (Geneletti, 2007) because they share the same identification formulas with natural effects. For the organic effects (Lok, 2015), note that they are identified as
$\operatorname{ODE}=\psi^{organic}(0,1)-\psi^{organic}(0,0),\operatorname{OIE}=\psi^{organic}(1,1)-\psi^{organic}(0,1),$ (22)
where
$\psi^{organic}(a_{1},a_{2})=\int_{w}\int_{z}\int_{m}\mathsf{E}[Y\mid A=a_{1},M=m,Z=z,W=w]\,\mathrm{d}\mathsf{P}(m\mid a_{2},z,w)\,\mathrm{d}\mathsf{P}(z\mid a_{1},w)\,\mathrm{d}\mathsf{P}(w),$
the equations $\operatorname{ODE}=\psi^{N}(0,1,0)-\psi^{N}(0,0,0)$ and $\operatorname{OIE}=\psi^{N}(1,1,1)-\psi^{N}(0,1,0)$ became obvious after we noticed that $\psi^{organic}(a_{1},a_{2})=\psi^{N}(a_{1},a_{2},a_{1})$.
For the randomized interventional effects (VanderWeele et al., 2014), we define
$\psi^{randomized}(a_{1},a_{2})=\int_{w}\int_{m}\int_{z}\mathsf{E}(Y\mid A=a_{1},z,m,w)\,\mathrm{d}\mathsf{P}(z\mid A=a_{1},w)\,\mathrm{d}\mathsf{P}(m\mid A=a_{2},w)\,\mathrm{d}\mathsf{P}(w),$
then $\operatorname{RIIE}=\psi^{randomized}(1,1)-\psi^{randomized}(1,0)$ and $\operatorname{RIDE}=\psi^{randomized}(1,0)-\psi^{randomized}(0,0)$. Note that
$\psi^{randomized}(a_{1},a_{2})$ $=\int_{w}\int_{m}\int_{z}\mathsf{E}(Y\mid A=a_{1},z,m,w)\,\mathrm{d}\mathsf{P}(z\mid A=a_{1},w)\,\mathrm{d}\mathsf{P}(m\mid A=a_{2},w)\,\mathrm{d}\mathsf{P}(w)$
$=\int_{w}\int_{z}\int_{m}\int_{z^{\prime}}\mathsf{E}(Y\mid a_{1},Z=z^{\prime},m,w)\,\mathrm{d}\mathsf{P}(z^{\prime}\mid a_{1},w)\,\mathrm{d}\mathsf{P}(m\mid a_{2},Z=z,w)\,\mathrm{d}\mathsf{P}(z\mid a_{2},w)\,\mathrm{d}\mathsf{P}(w)$
$=\int_{w}\int_{z}\int_{m}\int_{z^{\prime}}\mathsf{E}(Y\mid a_{1},Z^{\pi}=z^{\prime},m,w)\,\mathrm{d}\mathsf{P}(z^{\prime}\mid a_{1},w)\,\mathrm{d}\mathsf{P}(m\mid a_{2},Z=z,w)\,\mathrm{d}\mathsf{P}(z\mid a_{2},w)\,\mathrm{d}\mathsf{P}(w)$
$=\int_{w}\int_{z}\int_{m}\int_{z^{\prime}}\mathsf{E}(Y\mid a_{1},Z^{\pi}=z^{\prime},m,w)\,\mathrm{d}\mathsf{P}(z^{\prime}\mid a_{1},m,w)\,\mathrm{d}\mathsf{P}(m\mid a_{2},Z=z,w)\,\mathrm{d}\mathsf{P}(z\mid a_{2},w)\,\mathrm{d}\mathsf{P}(w)$
$=\mathsf{E}\Bigg{[}\mathsf{E}\bigg{[}\mathsf{E}\Big{[}\mathsf{E}\big{[}\mathsf{E}(Y\mid a_{1},Z^{\pi},M,W)\,\big{|}\,A=a_{1},M,W\big{]}\,\Big{|}\,A=a_{2},Z,W\Big{]}\,\bigg{|}\,A=a_{2},W\bigg{]}\Bigg{]},$
$=\psi^{R}(a_{1},a_{1},a_{2},a_{2}),$
where the third and fourth equalities hold because $Z^{\pi}\mid A,W\sim Z\mid A,W$ and $Z^{\pi}\mbox{$\perp\!\!\!\perp$}M\mid A,W$, respectively. This shows $\operatorname{RIDE}=\psi^{R}(1,1,0,0)-\psi^{R}(0,0,0,0)$ and $\operatorname{RIIE}=\psi^{R}(1,1,1,1)-\psi^{R}(1,1,0,0)$.
For the path-specific effects based on recanting twins (Díaz, 2024), we note that $\operatorname{RT}_{1}=E(Y_{S_{0}}-Y_{S_{1}})$, where
$E(Y_{S_{0}})$ $=\int_{w}E(Y\mid A=1,w)\,\mathrm{d}\mathsf{P}(w)$
$=\int_{w}\int_{z}\int_{m}E(Y\mid A=1,z,m,w)\,\mathrm{d}\mathsf{P}(m\mid A=1,z,w)\,\mathrm{d}\mathsf{P}(z\mid A=1,w)\,\mathrm{d}\mathsf{P}(w)$
$=\psi^{N}(1,1,1),$
and
$E(Y_{S_{1}})$ $=\int_{w}\int_{z}\int_{m}E(Y\mid A=0,z,m,w)\,\mathrm{d}\mathsf{P}(z,m\mid A=1,w)\,\mathrm{d}\mathsf{P}(w)$
$=\int_{w}\int_{z}\int_{m}E(Y\mid A=0,z,m,w)\,\mathrm{d}\mathsf{P}(m\mid A=1,z,w)\,\mathrm{d}\mathsf{P}(z\mid A=1,w)\,\mathrm{d}\mathsf{P}(w)$
$=\psi^{N}(0,1,1).$
For $\operatorname{RT}_{2}$, we note that $\operatorname{RT}_{2}=E(Y_{S_{1}}^{\prime}-Y_{S_{2}}^{\prime})$, where
$E(Y_{S_{1}}^{\prime})$ $=\int_{w}\int_{z}\int_{m}E(Y\mid A=0,z,m,w)\,\mathrm{d}\mathsf{P}(m\mid A=1,w)\,\mathrm{d}\mathsf{P}(z\mid A=1,w)\,\mathrm{d}\mathsf{P}(w)$
$=\int_{w}\int_{z}\int_{m}\int_{z^{\prime}}\mathsf{E}(Y\mid 0,Z^{\pi}=z^{\prime},m,w)\,\mathrm{d}\mathsf{P}(z^{\prime}\mid 1,m,w)\,\mathrm{d}\mathsf{P}(m\mid 1,Z=z,w)\,\mathrm{d}\mathsf{P}(z\mid 1,w)\,\mathrm{d}\mathsf{P}(w)$
$=\psi^{R}(0,1,1,1),$
and
$E(Y_{S_{2}}^{\prime})$ $=\int_{w}\int_{z}\int_{m}E(Y\mid A=0,z,m,w)\,\mathrm{d}\mathsf{P}(m\mid A=1,w)\,\mathrm{d}\mathsf{P}(z\mid A=0,w)\,\mathrm{d}\mathsf{P}(w)$
$=\int_{w}\int_{z}\int_{m}\int_{z^{\prime}}\mathsf{E}(Y\mid 0,Z^{\pi}=z^{\prime},m,w)\,\mathrm{d}\mathsf{P}(z^{\prime}\mid 0,m,w)\,\mathrm{d}\mathsf{P}(m\mid 1,Z=z,w)\,\mathrm{d}\mathsf{P}(z\mid 1,w)\,\mathrm{d}\mathsf{P}(w)$
$=\psi^{R}(0,0,1,1).$
For $\operatorname{RT}_{3}$, we note that $\operatorname{RT}_{3}=E(Y_{S_{2}}^{\prime\prime}-Y_{S_{3}}^{\prime\prime})$, where
$E(Y_{S_{2}}^{\prime\prime})=E(Y_{S_{2}}^{\prime})=\psi^{R}(0,0,1,1),$
and
$E(Y_{S_{3}}^{\prime\prime})$ $=\int_{w}\int_{z}\int_{m}\int_{z^{\prime}}E(Y\mid A=0,z,m,w)\,\mathrm{d}\mathsf{P}(m\mid A=1,z^{\prime},w)\,\mathrm{d}\mathsf{P}(z\mid A=0,w)\mathsf{P}(z^{\prime}\mid A=0,w)\,\mathrm{d}\mathsf{P}(w)$
$=\int_{w}\int_{z}\int_{m}\int_{z^{\prime}}\mathsf{E}(Y\mid 0,Z^{\pi}=z^{\prime},m,w)\,\mathrm{d}\mathsf{P}(z^{\prime}\mid 0,m,w)\,\mathrm{d}\mathsf{P}(m\mid 1,Z=z,w)\,\mathrm{d}\mathsf{P}(z\mid 0,w)\,\mathrm{d}\mathsf{P}(w)$
$=\psi^{R}(0,0,1,0).$
At last, for $\operatorname{RT}_{4}$, we note that $\operatorname{RT}_{4}=E(Y_{S_{3}}-Y_{S_{4}})$, where
$E(Y_{S_{3}})$ $=\int_{w}\int_{z}\int_{m}E(Y\mid A=0,z,m,w)\,\mathrm{d}\mathsf{P}(m\mid A=1,z,w)\,\mathrm{d}\mathsf{P}(z\mid A=0,w)\,\mathrm{d}\mathsf{P}(w)$
$=\psi^{N}(0,1,0),$
and
$E(Y_{S_{4}})$ $=\int_{w}E(Y\mid A=0,w)\,\mathrm{d}\mathsf{P}(w)$
$=\int_{w}\int_{z}\int_{m}E(Y\mid A=0,z,m,w)\,\mathrm{d}\mathsf{P}(m\mid A=0,z,w)\,\mathrm{d}\mathsf{P}(z\mid A=0,w)\,\mathrm{d}\mathsf{P}(w)$
$=\psi^{N}(0,0,0).$
Therefore, we show that
We skip the path-specific effects based on separable effects (Vo et al., 2024) as they share the same identification formulas with the path-specific effects based on recanting twins.
Appendix D Proofs of Theorem 1